# MMR-Mamba: Multi-Contrast MRI Reconstruction with Mamba and Spatial-Frequency Information Fusion

### v1

@@ IV-D1 Effectiveness of the Proposed Modules
To verify the validity of the proposed modules, we conducted ablation experiments on the BraTS dataset under 4$\times$ and 8$\times$ acceleration. The different configurations tested are denoted as follows: (a) a baseline model with all proposed modules removed; (b) the baseline model enhanced with the TCM; (c) the baseline model enhanced with the SFF; (d) the model augmented with both the TCM and SFF; and (e) the complete model incorporating all proposed modules. The results of these experiments are reported in Table II. As shown, incorporating the TCM improves the PSNR from 38.45 dB to 39.05 dB under 4$\times$ acceleration. while incorporating the SFF improves the PSNR to 40.49 dB under 4$\times$ acceleration Combining both TCM and SFF modules further boosts PSNR to 40.66 dB. Finally, the inclusion of the ASFF led to additional improvements. Similar trends were observed under 8$\times$ acceleration. Visual results on the BraTS dataset, including error maps, are shown in Fig. 5. As shown in the figure, the TCM module alone reduces the overall error, yet some structural details remain unrecovered, as indicated in the yellow box. Conversely, the SFF module effectively restores structural details but exhibits larger errors in the lateral ventricle areas, highlighted in red box. When both the TCM and SFF modules are employed, most regions and structures are adequately restored. Ultimately, integrating all proposed modules yields the best results. The quantitative and visualization results from our ablation experiments verify the effectiveness of each proposed module in enhancing the performance of multi-contrast MRI reconstruction.
@@ IV-D2 Analysis on Spatial Domain Fusion
To evaluate the effectiveness of our proposed spatial domain fusion module, we design three ablation experiments: (1) fusion through element-wise addition, (2) fusion through cross-attention, and (3) fusion through our proposed TCM module, donated as ‘Sum’, ‘CA’ and ‘TCM’ in Table III. The results indicate that element-wise addition of features from two modalities leads to the worst results. In contrast, our TCM module achieves superior performance, outperforming ‘Sum’ and ‘CA’ under both 4$\times$ and 8$\times$ acceleration, notably with minimal parameters.
@@
To verify the effectiveness of the proposed selective frequency fusion module, we design the following experiments: (1) fusion through element-wise addition and (2) fusion through our proposed SFF module, donated as ‘Sum’ and ‘SFF’ in Table IV, respectively. The results presented in the table show that element-wise addition of features from two modalities results in suboptimal performance. In contrast, our proposed SFF module significantly enhances performance, achieving a 0.43 dB improvement in PSNR under 4$\times$ acceleration and a 0.21 dB improvement under 8$\times$ acceleration.
@@ V Conclusion
This study explores the comprehensive and efficient integration of complementary information across modalities for multi-contrast MRI reconstruction. We present our MMR-Mamba framework, which integrates information through the TCM in the spatial domain and the SFF in the frequency domain, along with integrating the spatial-frequency features through the ASFF module. We conducted extensive experiments on the BraTS and fastMRI knee datasets, with results demonstrating the superiority of our proposed module in reconstructing MRI under different acceleration factors. This work contributes valuable insights for advancing research in multi-contrast MRI reconstruction, particularly in optimizing information fusion across spatial and frequency domains.

### v2

@@ IV-D1 Effectiveness of the Proposed Modules
To verify the validity of the proposed modules, we conducted ablation experiments on the BraTS dataset under 4$\times$ and 8$\times$ acceleration. The different configurations tested are denoted as follows: (a) a baseline model with all proposed modules removed; (b) the baseline model enhanced with the TCM; (c) the baseline model enhanced with the SFF; (d) the model augmented with both the TCM and SFF; and (e) the complete model incorporating all proposed modules. The results of these experiments are reported in Table II. As shown, incorporating the TCM improves the PSNR from 38.45 dB to 39.05 dB under 4$\times$ acceleration, while incorporating the SFF improves the PSNR to 40.49 dB under 4$\times$ acceleration. Combining both TCM and SFF modules further boosts PSNR to 40.66 dB. Finally, the inclusion of the ASFF led to additional improvements. Similar trends were observed under 8$\times$ acceleration. Visual results on the BraTS dataset, including error maps, are shown in Fig. 5. As shown in the figure, the TCM module alone reduces the overall error, yet some structural details remain unrecovered, as indicated in the yellow box. Conversely, the SFF module effectively restores structural details but exhibits larger errors in the lateral ventricle areas, highlighted in the red box. When both the TCM and SFF modules are employed, most regions and structures are adequately restored. Ultimately, integrating all proposed modules yields the best results. The quantitative and visualization results from our ablation experiments verify the effectiveness of each proposed module in enhancing the performance of multi-modal MRI reconstruction.
@@ IV-D2 Analysis on Spatial Domain Fusion
To evaluate the effectiveness of our proposed spatial domain fusion module, we design three ablation experiments: (1) fusion through element-wise addition, (2) fusion through cross-attention, and (3) fusion through our proposed TCM module, denoted as ‘Sum’, ‘CA’ and ‘TCM’ in Table III. The results indicate that element-wise addition of features from two modalities leads to the worst results. In contrast, our TCM module achieves superior performance, outperforming ‘Sum’ and ‘CA’ under both 4$\times$ and 8$\times$ acceleration, notably with minimal parameters.
@@
To verify the effectiveness of the proposed selective frequency fusion module, we design the following experiments: (1) fusion through element-wise addition and (2) fusion through our proposed SFF module, denoted as ‘Sum’ and ‘SFF’ in Table IV, respectively. The results presented in the table show that element-wise addition of features from two modalities results in suboptimal performance. In contrast, our proposed SFF module significantly enhances performance, achieving a 0.43 dB improvement in PSNR under 4$\times$ acceleration and a 0.21 dB improvement under 8$\times$ acceleration.
@@ V Conclusion
This study explores the comprehensive and efficient integration of complementary information across modalities for multi-modal MRI reconstruction. We present the MMR-Mamba framework, which effectively integrates information through the TCM module in the spatial domain and the SFF module in the frequency domain, along with integrating the spatial-frequency features through the ASFF module. In particular, the TCM module employs cross Mamba blocks to selectively supplement complementary information from the reference modality to the target modality, while the SFF module integrates global information in the Fourier domain and restores high-frequency signals for reconstructing structural details. We conducted extensive experiments on the BraTS and fastMRI knee datasets, demonstrating the superiority of our proposed framework in reconstructing MRI under different acceleration factors.